length of a plot is 57 meters and breadth is 47 meter.how many cent is the plot
Found 2 solutions by Alan3354, Cromlix:
Answer by Alan3354(69443)

What do you mean by "cent?"

Answer by Cromlix(4381)

length of a plot is 57 meters and breadth is 47 meter.
Area = length * breadth
Area = 57 * 47 = 2679 m^2
Perimeter = 2* length + 2* breadth
Perimeter = 2*57 + 2*47
Perimeter = 209 metres.
Hope this helps.
:-)